UT 20130920 AO#2 DX and AGw2 Calibration

Observer: DMiller, JChristou, JHill (LBTO)
Telescope Operator: GBechetti
SW Support: JKraus
AO Support: JChristou
Telescope Support: KNewton
Instrument Support: ESolheid

Summary

The good news: We have finished the pre-LUCI2 phase of the off-axis AGw2 alignment. That includes centering the guide probe; refining the guiding hotspot position; refining the pupil center position; adjusting M3 to minimize pupil wobble; and adjusting DX M2 to minimize binodal astigmatism.
The bad news: The sagging as a function of rotator angle has returned - Issue 4790. So while it is properly aligned optically, AGw2 is not performing well as a guider.

We lost some time to humidity at the beginning and the end of the night, but conditions were decent in between.

We verified that the residual pupil wobble for AGW2 (big W) is with the range of the camera lens.

Details

AGw2 Calibration

Open

02:00 Open chamber

02:15 Test of WFS communication to AdSec: Success

02:10 Geno set pointing

02:20 Doug S will do some LBC on the LEFT side. Geno authorize IRTC@RFBG and LBC@LeftPrime.

02:25 Put +100 Tilt on M2 and -100 CA to correct binodal. Start

DX M2 Manual Pointing Offsets are X=-5.107 RY=-982.

02:27 Send RIGHT preset but failed because too far from LLBC pointing position

02:29 irc_binocularcontrol, 'CLEARPRESET', /LEFT and /RIGHT

02:30 Left preset accepted

Close

02:33 94% Humidity frown, sad smile

Open

03:20 Reopening

Collimating

See Issue 4810

Setting the new guiding hotspot for AGw2

03:40 Used output of gcs_wfsc_display (x,y average removed) and command line % getHotspot right and % setHotSpot right -x 280 -y 225 to determine the center of the pinhole (i.e. the correct guiding hotspot)

03:50 Updated the GCS config file /home/telescope/TCS/Configuration/GCS/LUCI_IRTC_R.cfg with the
new guide hotspot of: X=280, Y=225 unbinned pixels.

03:56 Set "pseudomono" on right -- had been pure binocular before this.

Centering the rotator axis to hotspot

/home/lbto/data/20130921/center_agw2/

04:03 ACTIVE preset on right (LBC on left is passive)

04:05 image 0002 had peculiar hotspot (used for centering on IRTC last night).

04:09 Set RIGHTPOINTINGORIGIN to 0.0 0.0 to so the guide probe is put on axis on a preset.

restart PCS

04:10 ACTIVE preset on right (LBC on left is passive)

stop guiding

take test image (the one that always fails after you stop guiding)

Move guide probe to the nominal center -- setxy -u 2 -x 0.0 -y 612.5

04:1x readGuideCam right -e 80000 -- right_guider 0003 had a loopy circle ???? See JPEG attached below.

(JMH: 4 hours later the light bulb goes on, and we understand that the funny shape in our circle is caused by something loose in AGw2. This is the same thing that causes the pupil wobble. This indicates that it is something that affects both guider and wfs (e.g. the theta stage).

  • Circle for measuring rotator center with distortion:
    right_guider000003.jpg

04:18 readGuideCam right -e 80000 --
right_guider 0004 134,139 ---> 142,146 --->284,292
.............................IMAGE...........CCD.............UNBINNED

need -67 unbinned pixels in Y ---> +2000 microns RAD ooffset

04:37 Change AGw2 RAD ooffset from 417147 to 419147 in oacontrol.conf

Stop GCSR, stopAGw, stop and restart oacserver, startAGw, start GCSR

Lost the star

04:47 ACTIVE preset -- but didn't find star ????? WHY?

.............see the Night Log

Found star at IE=-6 CA=-131

Reapply Tilt=+100 to M2 and adjust CA by -100 to -231

05:25 ACTIVE preset to new BS star BS9186

05:33 readGuideCam right -e 90000 -- right_guider 0005 very jumpy

ptincrement 1 CA 5

05:37 readGuideCam right -e 90000 -- right_guider 0006

ptincrement 1 CA 10

05:40 readGuideCam right -e 90000 -- right_guider 0007

center of rotation is approx 130,107 ---> 138,115 ---> 276, 230 (matches the hotspot within the large measurement uncertainty -- the open-loop tracking is very jumpy even though the seeing is not so bad)

Refine AGw2 Pupil Wobble

/home/lbto/data/20130921/pupil_wobble_1/

05:47 ACTIVE preset BS9186 at EL=56 - but rotator doesn't want to move ?? Geno restarts rotator, and preset fails with jumpy acquisition offset.

05:53 Resend ACTIVE preset BS9186 - fails with jumpy acquisition offset again.

Preset misses the acquisition offset. Is this because of PM mode? See Issue 4811.

05:55 Authorize IRTC on right None on left.

05:56 ACTIVE preset BS9186 -- preset succeeds with smooth acquisition offset.

Doug is running pupil_collect. DX M3 Global offsets are zero.

06:28 DX M3 Global Offset Tip=+200, Tilt remains 0.

06:31 DX M3 Global Offset Tip=-50, Tilt=+200.

06:35 DX M3 Global Offset Tip=-50, Tilt=-200.

06:39 DX M3 Global Offset Tip=-150, Tilt=-500.

06:42 DX M3 Global Offset Tip=-300, Tilt=-500.

06:45 DX M3 Global Offset Tip=-300, Tilt=-700.

startWFS right (from the command line)

stopWFSing (from the GCSGUI)

06:51 pupil_collect

07:06 DX M3 Global Offset Tip=-150, Tilt=-350 . Absolute Tip=-558 Tilt=+330 RZ=-25.830
JMH: These were merged into oss.conf the following afternoon.

Stop here, even though the pupil wobble has converged to a line. We reopened Issue 4790.

Set New Pupil Center Position

New value in LUCI_IRTC_R.cfg wfscam_hotspot_x = 369.5 wfscam_hotspot_y=258.4 pixels

Field Data BS9107

/home/lbto/data/20130921/field/

07:45 ACTIVE preset to BS9107 at EL=52

07:47 ACQUIRE preset -- IE=+2 CA=-218

field_collect

GStar wfsc #
0 96-97
23 98-100
0 101
28 102-104
0 105
36 106-108
0 109-114
33 115-117
0 118
18 119-121
GStar wfsc #
0 122-124
39 125-127
0 128-129
24 130-132
0 133

Seeing on DIMM is ~0.9, Seeing on guider and wfs is ~0.7.

08:30 *Apply Tip-tilt correction for binodal to DX M2 Tip=+15 Tilt=+90 arcsec (absolute)
deltaIE=-15 ---> IE=-13 deltaCA=+10 ---> CA=-208*

DX M2 Manual Pointing Offsets are X=-4.596 Y=-0.766 Z=0 RX=147.3 RY=-883.8
JMH: These have been merged into the DXSM collimation files the following afternoon.
On 20130921 UT, Doug/John measured and adjusted the binodal astigmatism for
RFBG.  These corrections have been merged with the global offsets
below to make the final collimation values.  This has been applied in
DXSMIRTCFCollimation.dat for IRTC@RFBG and for all the other right
side focal stations.  Global offsets should now appear to be zero.

DX M2 Manual Pointing Offsets X=-4.596 Y=-0.766 Z=0 RX=147.3 RY=-883.8
DX M2 Global Offsets are      X=-1.569 Y=-0.368 Z=0 RX=-340.0 RY=-55.0
DX M2 Collimation Offsets     X=+3.543 Y=-2.930 Z=0.899 RX=-322.0 RY=+722.0
DX M2 New Collimation Offsets X=-2.595 Y=-4.064 Z=0.899 RX=-514.7 RY=-216.8

JMH: The corresponding IE=-13 CA=-208 pointing corrections have been merged into all the DX Gregorian poinitng models on 20130924.

Second set of field data

/home/lbto/data/20130921/field_2/

08:38 field_collect

GStar wfsc #
0 152-153
23 154-156
0 157
28 158-160
0 161-163
36 164-166
0 167-169
33 170-172
0 176-175
18 178-180
0 184-183
39 185-187

09:00 Now with AzCam 5.1 and in monocular mode, the cycle time of the guide loop is the exposure time plus 1.1 sec of readout (0.6 sec) and analysis (0.1 sec) and other (0.4 sec) including moving telescope . Thus with 2000 ms guider exposures on AGw2 we are guiding at 0.32 Hz.

DX AO Observational Robustness Testing

Humidity is rising......

Pyramid Pupil Wobble

09:21 ADAPTIVE_ACE preset to AO262 at EL=33 -- astigmatism LUT is off -- gs9 is a binary

09:28 RunAO

09:33 Offset -- failed -- did Bayside stages move?

09:36 ADAPTIVE_ACE preset to AO262 at EL=35

09:39 RunAO

09:50 acq sky_irtc_psf

Close

10:21 close for rising humidity

-- DougMiller - 20 Sep 2013
Topic attachments
I Attachment Action Size Date Who Comment
right_guider000003.jpgjpg right_guider000003.jpg manage 48 K 21 Sep 2013 - 08:03 JohnHill Circle for measuring rotator center with distortion
Topic revision: r10 - 25 Sep 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback